OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

one second let me write the equation.

OpenStudy (alexandervonhumboldt2):

ok

OpenStudy (anonymous):

\[if f(x)=6x-4, evaluate f(x)/-4 for x=8\]

OpenStudy (alexandervonhumboldt2):

(6x-4)/-4 when x=8?

OpenStudy (alexandervonhumboldt2):

(48-4)/-4=44/-4=-11